Opened 6 years ago

Last modified 6 years ago

#416 assigned flight processing

BB11_01, flight day 244b/2011, Brecon Beacons

Reported by: adbe Owned by: mggr
Priority: alpha 4 high Milestone: 2011 data processing completion
Component: Processing: general Keywords:
Cc: mggr Other processors:

Description (last modified by knpa)

Data location: /users/rsg/arsf/arsf_data/2011/flight_data/uk/BB11_01-2011_244b_Brecon_Beacons

Data arrived from ARSF via network transfer on 02/09/2011

BB11_01 Scientific objective:

BB11_01 Priority: a4h

PI: Shaun Lewis (Brecon Beacons National Park Authority)

Sensors:

Camera (11/10/2011)
Eagle (23/11/2011)
Hawk (23/11/2011)
LiDAR (23/11/2011)

Special note

Atmospheric correction requested. Please check with Ben, Steve or Mike prior to creating a normal delivery.

Change History (21)

comment:1 Changed 6 years ago by adbe

Processed navigation

comment:2 Changed 6 years ago by jaho

Started camera processing. Removed image 94.

comment:3 Changed 6 years ago by jaho

Camera ready for delivery check.

comment:4 Changed 6 years ago by jaho

Lidar ready for delivery check.

comment:5 Changed 6 years ago by adbe

Hyperspectral being processed with NextMap DEM

comment:6 Changed 6 years ago by adbe

Created hyperspectral delivery.
Note:
*Used NextMap DEM
*Needs to be atmospherically corrected after check

comment:7 Changed 6 years ago by emca

Delivery checked Camera and LiDAR data.

Need to add the flight letter to the project details section of both readme's and to the KML file.
Also add note to clarify that this data is not atmospherically corrected (only the Hyperspectral data will be).

All data looks good and ready to delivery once minor changes are made.

comment:8 Changed 6 years ago by jaho

Camera and LiDAR sent to Shaun Lewis on hard drive 114 (11/10/2011).

comment:9 Changed 6 years ago by benj

ASD Fieldspec and Microtops measurements received from BBNPA and stored in the flight directory under ground_measurements (symlink to the 244a directory).

comment:10 Changed 6 years ago by knpa

Gary says that the PI should be able to get their own vectors via PGA.
Do not wait for vectors to make delivery but deliver with caveat that we have not made this usual check.

comment:11 Changed 6 years ago by jaho

Reprocessed LiDAR flightlines with pitch offset -0.00291419. Created new ASTER dem in place of the old SRTM one. Updated elevation differences. The data seems to look better. New delivery at: delivery/BB11_01-244b-lidar-20111117.

comment:12 Changed 6 years ago by knpa

Doing HS delivery check.

comment:13 Changed 6 years ago by knpa

Delivery check complete. All looks good apart from:

  1. File names should really be 1-8, not 8-15.
  2. Project information xml should have flight letter in it, if it is specific to 244b.
  3. Mention in Readme that we didn't check against vectors.
  4. Bad pixel files unix2dos

comment:14 Changed 6 years ago by adbe

LiDAR re-delivery check:

  • There are 2 files: u2dtmprvDzfc, u2dtmpUgzooG; in the ascii directory.
  • Mention vectors in ReadMe

Otherwise, ready to deliver

comment:15 Changed 6 years ago by adbe

Hyperspectral:

  • Renamed project xml
  • Mentioned vectors in ReadMe
  • Done unix2dos on .txt files

comment:16 Changed 6 years ago by jaho

Updated LiDAR ReadMe to mention vectors. Removed u2dtmprvDzfc and u2dtmpUgzooG.

comment:17 Changed 6 years ago by adbe

  • Description modified (diff)

Hyperspectral(No atmospheric correction) and LiDAR(correct pitch) delivered to Shaun Lewis on 23/11/2011 on disk 117

comment:18 Changed 6 years ago by mark1

Hyperspectral hawk bad pixel files and XML files have been re-created and re-delivered to PI (via email). This corrects sample number in bad pixel text files and Eagle IFOV value in XML and extra comments included.

comment:19 Changed 6 years ago by adbe

Hyperspectral sct's

lineeaglehawk
8-0.05-0.07
9-0.06-0.04|
10-0.05-0.06
11-0.02-0.04
12-0.05-0.06
13-0.01-0.03
14-0.06-0.08
15-0.03-0.05

comment:20 Changed 6 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Description modified (diff)
  • Owner changed from adbe to knpa
  • Status changed from new to assigned

comment:21 Changed 6 years ago by knpa

  • Component changed from Archiving to Processing: general
  • Owner changed from knpa to mggr

Taken back off archiving as some atmospheric stuff still to be done on this project I believe...

Note: See TracTickets for help on using tickets.